OpenStudy (anonymous):

Word problem: Sam has 30 nickels and dimes worth $2.40. How many nickels does he have? My problem: I am having trouble making this into an equation. I don't need help solving it, just making the equation. That's all

OpenStudy (anonymous):

In terms of n for nickles and d for dimes what would the equation look like for number of coins?

OpenStudy (anonymous):

I have no clue at the moment. In fact I'm having the worst brain fart ever.

OpenStudy (anonymous):

How many coins do you have?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

So n+d = 30 this would be your first equation as it represents the number of coins. Do you understand?

OpenStudy (anonymous):

Oh! yes, yes I do now! Now I remember what to do. Thanks

OpenStudy (anonymous):

How will you write the second equation?

OpenStudy (anonymous):

ummmm, d+2.40+d=30? Oh great, now i think I'm brain farting again. maybe

OpenStudy (anonymous):

I was just going by an example I have in my notes

OpenStudy (anonymous):

It has to equal $ 2.40

OpenStudy (anonymous):

How much is a dime worth and how much is a nickel worth?

OpenStudy (anonymous):

a nickle=5 cents and a dime= 10 cents

OpenStudy (anonymous):

Or .05 and .10.. So we need an equation that represents the number of nickels times .05 + the number of dimes times .10 = $ 2.40..so it will look like what?

OpenStudy (anonymous):

.05n+.10d=2.40

OpenStudy (anonymous):

Excellent. To solve you will use substitution

OpenStudy (anonymous):

Okay, so then I would change it to look like .05n+.10(n+30)=2.40 which will be n=$19.84. I think. I'm pretty sure